Cara used 4 1/6 sheets of construction paper to make 1/6 of her poster board project. At that rate, how many sheets will she use for the whole project?
Answer by greenestamps(13198)



6 times as much as she used for the first 1/6 of the project. 6*4 = 24; 6*1/6 = 1; 24+1 = 25